OpenStudy (anonymous):

Find the prime factors of 19404 first.

OpenStudy (anonymous):

The Prime Factors of 19404: 2^2 • 3^2 • 7^2 • 11

OpenStudy (anonymous):

can u explain in brief plzz

OpenStudy (anonymous):

@EswarVarma Sorry, I'm a little busy but this link leads to a very helpful video on prime factorization if you are interested! https://www.khanacademy.org/math/pre-algebra/factors-multiples/prime_factorization/v/prime-factorization

OpenStudy (anonymous):

kk thank u

OpenStudy (anonymous):

19404/2=9702 9702/2=4851 4851/3=1617 1617/3=539 539/7=77 77/7=11 2 2 3 3 7 7 11